### v1

# Infinite magneto-resistance and bipolar effect in spin valves driven by spin batteries

Abstract
It is shown that spin valves under suitable symmetry conditions exhibit an ON-OFF response to a spin battery. While a spin valve driven by a charge battery displays the usual GMR (Giant Magneto-Resistance), a pure spin current or pure spin accumulation can generate an infinite magneto-resistance effect(IMR). In practice the effect is curtailed by asymmetry but the achievable magneto-resistance ratio is still predicted to be unusually large in several example setups. It is closely related to the bipolar effect (BE) of Johnson transistor which can be triggered under the same symmetry conditions.
1 Introduction
This paper introduces a new magneto-resistive effect, the infinite magneto-resistance (IMR) related to the giant magneto-resistance spin valve (GMR [1, 2]) and to the bipolar spin transistor (BST) of Johnson [3, 4].
The IMR does not rely on specific materials such as half-metals to increase GMR (Heusler alloys as in [5]) or coherent tunneling through MgO barriers [6, 7]. It consists in achieving a zero signal for a system with 2 magnetically active parts (like two ferromagnetic layers in a spin valve) so that the optimistic (resp. pessimistic) MR ratio tends to infinity (resp. unity), hence the naming of the effect as IMR. Its main characteristics are: (1) the use of a spin battery instead of a charge battery; (2) the reliance on symmetry to get rid of the offsets preventing a bipolar response in Johnson BST; (3) the reliance on Johnson-Silsbee charge-spin coupling [8] to measure the effect. Of course, the effect is curtailed by asymmetries which always exist so that the MR ratio can never be infinite and it will be one of the objectives of the paper to quantify to impact of slight asymmetries and show that the MR ratio can still be very large.
The starting point of our proposed effect is the BST. In the latter, one has a spin valve like structure $F1-N-F2$ with two collinear ferromagnets $F1$ and $F2$ connected by a paramagnetic metal $N$. A charge current flows through the first ferromagnet $F1$ but NOT through the second $F2$. Thanks to spin injection it induces a spin current and spin accumulation which spill into the second ferromagnet $F2$. The spin accumulation in $F2$ is detected as a charge voltage through Johnson-Silsbee charge spin coupling. It was predicted that the signal would be bipolar as a function of the magnetization polarization (up or down): $V_{c}\propto s\mathrm{gn}\left(M_{z}\right)\propto\pm V_{1}$. Experimentally the bipolarity has proved elusive to detect and one observes instead
$V_{c}=V_{0}+V_{1}\,\sigma$ (1)
(where $\sigma=\pm$) with a voltage offset $V_{0}$ [9, 10, 3, 11, 12, 13, 14, 15, 16]. The origin of these offsets have been credited to various causes such as current inhomogeneities [11], heating [14, 15, 16], departures from one dimensional geometry [9].
Conceptually one can deconstruct the BST into three components: (1) a spin battery, made up of the ferromagnet $F1$ when a charge current flows through it; (2) a load $F2$ for the spin current and spin accumulation created by the spin battery; (3) a charge (voltage or current) measurement apparatus.
The systems we consider will be driven by a spin battery like the original BST but differ in that the load is not a single ferromagnet but has at least two magnetically active regions. For instance for collinear systems, we will have in general a voltage response
$V_{c}=V_{0}+V_{1}\,\sigma_{1}+V_{2}\,\sigma_{2}+V_{12}\,\sigma_{1}\,\sigma_{2}$ (2)
where $\sigma_{1/2}=\pm$ track the magnetization directions. The voltage offsets come from $V_{0}$ and $V_{12}$. But should they vanish due to some some symmetry, one would then have
$V_{c}=V_{1}\,\sigma_{1}+V_{2}\,\sigma_{2}$ (3)
exhibiting a bipolar response when both magnetizations are switched ($\sigma_{1/2}\longrightarrow-\sigma_{1/2}$). Suppose now that due to a stronger symmetry $V_{1}=V_{2}$ (resp. $V_{1}=-V_{2}$); $V_{c}$ will then obviously vanish whenever the two magnetizations are antiparallel (resp. parallel). This implies that the optimistic MR ratio $\left(V_{HIGH}-V_{LOW}\right)/V_{LOW}$ will be infinite. This is the IMR which the main topic of this paper. In such a spin valve, a single magnetization needs to be switched; if instead both are switched, one recovers a BE.
Aside from symmetry, the other main ingredient is the use of a spin battery instead of a regular charge battery. The reason is that oftentimes the response $r_{c}\left(\sigma\right)=\partial V_{c}/I_{c}$ is even in $\sigma$ (the magnetization direction), $r_{c}\left(-\sigma\right)=r_{c}\left(\sigma\right)$ which implies that $V_{1}=0$ so that the IMR is not possible with a pure charge battery (since it lacks any magnetically active region). In contrast a spin battery will induce an odd term in the voltage response, which is required for the BE and IMR.
The third ingredient is the reliance on Johnson-Silsbee charge-spin coupling (the coupling of charge and spin transport in magnetic systems) to get a (charge) voltage or current reading from magnetic changes. The advantage of a voltage measurement over a current one is that it avoids Joule heating. But theoretically either one could be done.
What is of paramount importance is to have symmetrical setups which will allow to get rid by construction of any offsets. A simple recipe is to have two copies of the same system as in Fig. 1-a: the offsets obviously cancel out. This is an example of what we will christen as ’twin systems’.
Figure 1: (a) The voltage measured across two copies of the same system (and batteries) vanishes by symmetry. (b) FNF valve connected to a spin battery and a voltmeter. If the combined system is symmetrical under left-right exchange, then the voltage must vanish by symmetry. The layers can be viewed as parallel to the figure plane (as in non-local geometries) or to be perpendicular. SB: spin battery. (c) Same as (b) but depicted with an FNF nanopillar playing the role of the SB. (d) CPP trilayer driven symmetrically by two spin batteries. For parallel magnetizations, the voltage again vanishes (for symmetric probes). The black arrows depict the magnetization directions.
Another example of a ’twin system’ is given in Fig. 1-b. It is an $FNF$ valve in a non-local geometry with identical ferromagnets and layers parallel to the figure plane; it could also be considered to be a trilayer with layers perpendicular to the figure plane. A spin battery injects a spin current through the intermediate $N$ layer. The charge voltage across the trilayer obviously vanishes when the ferromagnets have parallel magnetizations but is non-zero for anti-parallel ones. In practice the vanishing of the signal is constrained by the degree of symmetry. Suppose for instance that the widths of the ferromagnets are equal to $l_{m}\pm\delta l$. Then we expect a scaling:
$\frac{V_{c}^{AP}-V_{c}^{P}}{V_{c}^{P}}\sim\frac{l_{m}}{\delta l}$ (4)
so for widths equal with an accuracy of $1\%$, the ratio reaches $10^{2}=10^{4}\>\%$. Other geometrical asymmetries will of course affect this rough estimate. For the same twin setup, the BE is achieved by starting from an antiparallel configuration and flipping both magnetizations.
There is a wide variety of spin batteries currently in use and among others: spin injection devices (a charge current gets polarized by traversing a ferromagnet [17, 18, 19]); spin Hall devices (a transverse spin current is generated when there is a longitudinal charge current due to spin-orbit interactions [20]); spin pumps (a rotating bulk magnetization engenders a spin current, as in the ferromagnetic resonance battery [21, 22]); tunnel junctions relying on photoexcitation (circularly polarized light at FM/semiconductor tunnel junctions excites spin polarized carriers [23]); devices relying on the spin Seebeck effect (a thermal gradient can generate a spin current) [24, 25]. Fig. 1-c shows the twin system from Fig. 1-b with an FNF nanopillar acting as a spin battery.
There are other possible setups than ’twin setups’ which will exhibit IMR and bipolarity, as a result of symmetry. For instance consider a CPP symmetric trilayer where instead of a charge battery, one connects the layers to two spin batteries as in Fig. 1-d. (The CPP qualifier, current perpendicular to plane, comes from the fact that if one replaced the voltage probes by current ones, the charge current would be perpendicular to the layers.) This setup exhibits IMR for identical incoming spin currents and parallel magnetizations. We will discuss later a similar setup with opposite spin currents; one then has IMR for antiparallel magnetizations.
The latter case is actually easy to understand in terms of a resistor analogy. We consider a spin accumulation source for the spin battery (a spin current source will yield similar results); in this example we will measure a charge current. In the resistor analogy for a CPP symmetric spin valve with anti-parallel magnetizations driven by a charge battery, currents flowing in each spin channel are equal $I_{\uparrow}=I_{\downarrow}$. The key observation is to ask what would happen if the spin channels had a reverse bias (opposite voltages) as in Fig. 2-(b) : the currents in each spin channel are now opposite $I_{\uparrow}=-I_{\downarrow}$; obviously no charge current can thus flow out of the device (since $I_{c}=I_{\uparrow}+I_{\downarrow}$) in contrast to Fig.2-(a) (parallel case). But a reverse bias is precisely what a spin accumulation source can achieve: indeed a spin battery splits the chemical potentials for each spin channel $V_{\uparrow}\neq V_{\downarrow}$ and additionally $V_{\uparrow}=-V_{\downarrow}$ since $V_{c}=0$ in the absence of a charge voltage applied to the system. Replacing a charge battery with a spin battery therefore generates reversed biases and opposite currents in each spin channel (for a symmetric spin valve), so that in the anti-parallel state, there is no outgoing charge current: the corresponding ratio $\left(I_{c}^{AP}-I_{c}^{P}\right)/I_{c}^{AP}$ goes to infinity (IMR effect).
Figure 2: In the resistor network model, symmetric CPP spin valve driven by a spin battery which injects a spin current and induces spin accumulation. The latter is equivalent to reverse biasing the two spin channels. The currents (in electrical units) in each spin channel are $I_{\uparrow}$ and $I_{\downarrow}$. (a) Parallel magnetizations for ferromagnets $F1$ and $F2$ : non-vanishing charge current $I_{c}^{P}\neq 0$. (b) Anti-parallel magnetizations: current $I_{c}$ at detector is vanishing since the currents in each channel are opposite. Therefore one has an infinite resistance state.
2 Response function, BE & IMR.
2.1 Trans-resistance.
Let us define a charge-spin trans-resistance encoding the response to a spin battery:
$r_{cs}=\left(\frac{\partial V_{c}}{\partial I_{s}}\right)_{I_{c}=0}.$ (5)
It has the dimension of a resistance. The basic geometries we consider are therefore 3 terminal ones, two terminals used for the voltage measurement and one terminal connected to the spin battery. We will make the hypothesis that the spin battery is a perfect spin current source which outputs a constant spin current $I_{s}$, whatever the load it is attached to. This is not a very stringent condition. It is easy to take into account spin backflow into the battery through the internal spin resistance concept [26], which in practice will translate into a reduced response. Let us suppose that the load comprises $N=2$ magnetically active regions, whose magnetizations are collinear (but maybe have different directions) and can be controlled independently. One can write therefore: $r_{cs}=r_{cs}\left(\sigma_{1},\;\sigma_{2}\right)$ where $\left\{\sigma_{1}=\pm 1,\;\sigma_{2}=\pm 1\right\}$ index the magnetizations of the $2$ regions relative to a given quantization axis. $r_{cs}$ is simply a multilinear polynomial in $\sigma_{1}$ and $\sigma_{2}$:
$r_{cs}\left(\sigma_{1},\,\sigma_{2}\right)=r_{0}+r_{1}\,\sigma_{1}+r_{2}\,\sigma_{2}+r_{12}\,\sigma_{1}\,\sigma_{2}.$ (6)
The resistance $r_{cs}$ can take four independant values.
2.2 Bipolar effect (BE).
This requires the concurrent switching of both magnetizations. One has a bipolar effect if for some values of $\sigma_{1}$ and $\sigma_{2}$
$r_{cs}\left(-\sigma_{1},-\sigma_{2}\right)=-r_{cs}\left(\sigma_{1},\sigma_{2}\right);$ (7)
which implies that $r_{0}=-r_{12}\,\sigma_{1}\,\sigma_{2}.$ In the special case $r_{0}=0,\;r_{12}=0,$ the bipolarity will occur for any values of $\sigma_{1}$ and $\sigma_{2}$. The twin system of Fig.1-a obeys
$r_{cs}\left(\sigma_{2},\sigma_{1}\right)=-r_{cs}\left(\sigma_{1},\sigma_{2}\right)$ (8)
so that $r_{0}=0,\;r_{12}=0,$$r_{1}=-r_{2}$. For Fig. 1-b, eq.(8) is also true since a symmetry through plane $(\mathcal{P})$ parallel to the layers and perpendicular to the Figure maps each ferromagnetic layer unto the other but leaves the spin battery invariant. But the voltage probes are exchanged across that plane so that $V_{c}\longrightarrow-V_{c}$. It follows that: $r_{cs}\left(\sigma_{1},\,\sigma_{2}\right)=r_{1}\cdot\left(\sigma_{1}-\sigma_{2}\right).$ If one starts from an antiparallel configuration and swaps both magnetizations, one therefore has a bipolar response:
$r_{cs}\left(\sigma_{1},\,-\sigma_{1}\right)=2r_{1}\sigma_{1}=-r_{cs}\left(-\sigma_{1},\,\sigma_{1}\right).$ (9)
2.3 IMR.
For the twin system of Fig.1-a & b, it also follows from eq.(8) that $r_{cs}^{P}=r_{cs}\left(\sigma_{1},\,\sigma_{1}\right)=0$ while $r_{cs}^{AP}\neq 0$. We have a parallel IMR. The pessimistic MR ratio is 100% and the optimistic MR ratio is infinite. If one starts from an antiparallel setup and one switches a single magnetization, the response vanishes (IMR). A similar symmetry condition can not be enforced when there is a single $N=1$ magnetized region because in order to detect a charge voltage through Johnson-Silsbee charge-spin coupling, the measurement probes must be positioned asymmetrically.
Since the condition for IMR also allows the BE, let us compare the two effects for the setups of Fig. 1-a, 1-b. IMR is observed by switching a single magnetization while bipolarity requires switching two magnetizations, which may or may not be possible or desireable in a given setup. The amplitude of the IMR effect is $\left|r_{cs}^{AP}-r_{cs}^{P}\right|=2r_{1}$, while the bipolar effect has an amplitude which is twice as large. There is therefore a trade-off between the two effects: IMR is smaller but easier to observe, requiring a single magnetization to be switched, while bipolarity has the larger signal but requires both magnetizations to be switched.
Let us compare with the regular charge resistance $r_{c}=\left(\frac{\partial V_{c}}{\partial I_{c}}\right)$; the latter has a similar dependence in general. But if the system is invariant when both magnetizations are reversed, one has the simpler dependence: $r_{c}\left(\sigma_{1},\,\sigma_{2}\right)=r_{c0}+r_{c12}\,\sigma_{1}\,\sigma_{2}.$ One recovers the usual GMR for a spin valve subjected to a charge battery. An infinite MR ratio is much more difficult to engineer. For instance, let us consider Juilliere model for TMR. Then: $r_{c}\propto\left(1+P_{1}P_{2}\right)$ where $P_{1/2}$ is a DOS spin polarization. For identical layers, $\left|P_{1}\right|=\left|P_{2}\right|=P$. An infinite MR ratio would then require $\left(1-P^{2}\right)=0$, that is $P=1$ or a half-metal. This is a very stringent condition when compared with the IMR discussed above.
In GMR spin valves, one usually anchors one of the two ferromagnets so that a single ferromagnet can be switched under normal operating conditions. Doing so here would obviously prevent the symmetries enabling the IMR and bipolar effects. This is another difference with conventional GMR spin valves. The switching field needs to be tailored according to the effect one wishes to trigger, whether it be IMR or the bipolar effect.
3 Example setups.
3.1 Twin setup with shared spin battery.
To quantify the impact of asymmetries, one can make explicit calculations for the geometry of Fig.(1-b) in the one dimensional limit, using Valet-Fert drift diffusion model. We introduce $P_{F1}$ and $P_{F2}$ are the conductivity polarizations for ferromagnets $1$ $2$, contact resistance $r_{c}$, conductance polarization $P_{c1}$ and $P_{c2}$ at the interfaces between the ferromagnets $1$ 2 and the central paramagnetic layer, contact resistances $r_{c}^{\prime}$, conductance polarization $P_{c1}^{\prime}$ and $P_{c2}^{\prime}$ at the interfaces between the ferromagnets $1$ 2 and the measurement leads. We also define $l_{1/2}$ the ferromagnet layer widths relative to $l_{F}$ the spin relaxation length of the ferromagnets. Defining $l_{m}=\frac{l_{1}+l_{2}}{2},\;\delta l=\frac{l_{1}-l_{2}}{2}$ as well as $\sinh_{m}=\sinh l_{m},\;\cosh_{m}=\cosh l_{m},$
$X$ $=\frac{r_{N}+r^{\prime}_{c}}{r_{F}},\>f_{m}=\frac{\sinh_{m}+X\>\cosh_{m}}{\cosh_{m}+X\>\sinh_{m}},$ (10)
one finds after tedious calculations that $r_{cs}$ splits into 3 parts, a bulk contribution (I) and two interfacial ones (II & III):
$r_{cs}^{I}$ $=\left(P_{F1}-P_{F2}\right)\>\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right]\>\frac{r_{F}}{2}+\delta l\>\left(P_{F1}+P_{F2}\right)\>\frac{r_{F}}{2}\>\alpha,$ (11)
$r_{cs}^{II}$ $=\left(P_{c1}-P_{c2}\right)\>\frac{r_{c}}{2}+\delta l\>\left(P_{c1}+P_{c2}\right)\>\frac{r_{c}}{2}\>\beta,$ (12)
where $\alpha$ and $\beta$ are factors which depend on the geometry and the various spin resistances (for their exact expressions, we refer to [27]).
$r_{cs}^{III}$ $=\left(P_{c1}^{\prime}-P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)+\delta l\>\left(P_{c1}^{\prime}+P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$ (13)
One can compute a resistance ratio as $MR=\frac{R_{AP}-R_{P}}{R_{P}}$ where the subscripts $P/AP$ refer to parallel or anti-parallel magnetizations; it follows that:
$R_{P}=\circ\left(\delta l\right),\qquad MR\propto\delta l^{-1}$ (14)
Asymmetries in transverse directions should in principle be taken into account so that most generally one expects
$MR\propto\left(\sum_{i=x,y,z}C_{i}\>\frac{\delta l_{i}}{l_{i}}\right)^{-1}$ (15)
but for thin films, the dominant contribution will come from the width so that other asymmetries can be neglected. As an illustration, consider a Co-Cu-Co trilayer. For all the following figures, we refer to [28]; Piraux et al find at $77\>K$ that $l_{F}=36\>nm$. MSU Reilly et al have data consistent with a large $l_{F}$ maybe larger than $60\>nm$ at $4.2\>K$. If the widths are about $10\pm 0.5\,nm$, $\delta l\sim\frac{1}{60}$ so up to numerical factors $MR\sim 6000\%.$
It is also important to have an idea of the range of the HIGH signal. Let’s use MSU results for CoCu at $4.2\>K$ and for simplicity assume that $P_{c}\sim P_{c}^{\prime}=0.77\pm 0.05$, $P_{F}=0.46\pm 0.05,$ $2r_{c}A\sim 2r_{c}^{\prime}A=1.02\pm 0.04\>f\Omega m^{2}$, $t_{Co}=2-10\,nm,\>l_{Co}>60\>nm$ and $\rho_{F}(Co)=75\pm 5\>n\Omega m,$ $\rho_{N}(Cu)=6\pm 1\>n\Omega m.$ Then $X\sim 1$, $f_{m}\sim 1$. So for AP ferromagnets:
$r_{cs}^{high}\,A\sim 1\>f\Omega m^{2}.$ (16)
This will be in the range of $R=1\>\Omega$ for $A=10^{-3}\mu m^{2}$ which corresponds to a nanopillar.
3.2 Twins with individual spin batteries.
We consider two identical systems connected to a spin battery and interconnected by voltage probes (Fig.(1-a)). This geometry is better suited to the observation of IMR than the previous one because it is easier to switch a single ferromagnet since they can be well separated spatially. One must take into account spin batteries asymmetry $\delta I_{s}$; one finds a MR ratio
$R_{P}=\circ\left(\delta l\right)+\circ\left(\delta I_{s}\right),\qquad MR\propto\left(\frac{\delta l}{l}+C\,\frac{\delta I_{s}}{I_{s}}\right)^{-1}.$ (17)
For very wide ferromagnets, it can be shown that it is better to have $X=\frac{r_{N}+r^{\prime}_{c}}{r_{F}}\sim 1$ so that $C\sim 1-2$. In the opposite limite of very thin ferromagnets, $C\propto X$, the MR ratio will be boosted by transparent contacts ($r_{c}^{\prime}\ll\left(r_{N},\;r_{F}\right)$) which lower $C$. For very resistive contacts such that $\left(r_{c}^{\prime},\>r_{c}\right)\gg\left(r_{N},\;r_{F}\right)$ and $X\gg 1$, one finds that $C\propto 1/X$ provided the ferromagnets are not too thin. The reason for this potential enhancement is that the resistive contacts hinder the backflow of spin accumulation and allow its build-up. But the spin current asymmetry will likely limit any such improvement. The HIGH signal is dominated by interfaces in the limit of very thin ferromagnets: $r_{cs}^{high}\longrightarrow 2\>\left\{P_{c}\,r_{c}+P_{c}^{\prime}\,r_{c}^{\prime}\right\}$ while for very thick layers: $r_{cs}^{high}\longrightarrow 2\>\left\{P_{c}\,r_{c}+P_{F}\,r_{F}\right\}.$ Using the same parameters as above, one finds again a range $r_{cs}^{high}\,A\sim 1\>f\Omega m^{2}.$
3.3 Symmetric CPP spin valves operated by two spin batteries. Parallel IMR.
One can define new transport coefficients per:
$V_{c}=r_{cs}^{L}\>I_{s,L}+r_{cs}^{R}\>I_{s,R}.$ (18)
Since there is a new control parameter, it is easy to get a zero response (IMR) by adjusting one spin current relative to the other. One can then sweep the value of, say, the spin current $I_{s,L}$ until $I_{s,L}=-\frac{r_{cs}^{R}\>I_{s,R}}{r_{cs}^{L}}.$ No symmetry is required but the trade-off is that (1) the system is larger since there are two spin batteries, (2) while an adjustment is easy to do in a lab context, it is much less convenient to do so in industrial applications.
It is therefore advantageous to engineer symmetrical systems which can have reproducible behaviours. Consider for instance Fig.1-d which shows a spin valve in a CPP geometry (if the voltmeter was shorted, a charge current would flow in a direction orthogonal to the layers). Note that the setup in Fig. 1-d could represent also a nano-pillar or a non-local geometry in addition to the usual multi-layer thin films. The spin current from the batteries are oriented to be incoming in the load (outgoing from the batteries). If the ferromagnetic layers and the spin batteries are identical, there is a plane of symmetry $\mathcal{P}$ in the paramagnetic layer parallel to the layers. Obviously $V_{c}=0$ when the magnetizations are parallel. One has a parallel IMR. If one starts with antiparallel magnetizations and switches both of them, one gets a bipolar effect.
After rotating the setup around an axis in the middle of the paramagnetic layer:
$V_{c}\longrightarrow-V_{c},$ $\>I_{s,L}\longrightarrow I_{s,R},\>\left(\sigma_{1},\>\sigma_{2}\right)\longrightarrow\left(\sigma_{2},\>\sigma_{1}\right).$ (19)
Therefore:
$V_{c}$ $=r_{cs}^{L}\left(\sigma_{1},\>\sigma_{2}\right)\>I_{s,L}+r_{cs}^{R}\left(\sigma_{1},\>\sigma_{2}\right)\>I_{s,R},$ (20)
$-V_{c}$ $=r_{cs}^{L}\left(\sigma_{2},\>\sigma_{1}\right)\>I_{s,R}+r_{cs}^{R}\left(\sigma_{2},\>\sigma_{1}\right)\>I_{s,L},$ (21)
which implies
$r_{cs}^{L}\left(\sigma_{1},\>\sigma_{2}\right)$ $=-r_{cs}^{R}\left(\sigma_{2},\>\sigma_{1}\right).$ (22)
Therefore for identical currents $I_{s,L}=I_{s,R}=I_{s}$
$V_{c}$ $=\left[r_{cs}^{L}\left(\sigma_{1},\>\sigma_{2}\right)-r_{cs}^{L}\left(\sigma_{2},\>\sigma_{1}\right)\right]\>I_{s}$ (23)
which is antisymmetric under exchange of the magnetizations, $V_{c}\left(\sigma_{2},\>\sigma_{1}\right)=-V_{c}\left(\sigma_{1},\>\sigma_{2}\right).$ This implies a parallel IMR or vanishing response for parallel magnetizations:
$V_{c}\left(\sigma_{1},\>\sigma_{1}\right)=-V_{c}\left(\sigma_{1},\>\sigma_{1}\right)=0$ (24)
and a bipolar effect for antiparallel magnetizations when flipping both of them for identical currents: $V_{c}\left(\sigma_{1},\>-\sigma_{1}\right)=-V_{c}\left(-\sigma_{1},\>\sigma_{1}\right).$ Detailed calculations on an FNF trilayer exhibiting such a parallel IMR are given elsewhere [27]. In the limit of infinite ferromagnets, one finds
$MR=\frac{V_{c}^{AP}-V_{c}^{P}}{V_{c}^{P}}\sim\frac{I_{s,0}^{m}}{\delta I_{s}}$ (25)
3.4 Anti-parallel IMR in CPP MTJ & FNF trilayer.
It is also possible to engineer an anti-parallel IMR. Fig.3-a shows such a setup. A CPP trilayer is connected symmetrically to two spin batteries which inject opposite spin currents. If the voltage probes are also symmetrical, the charge voltage must vanish as shown in Fig.3. Indeed let us rotate the setup by an angle $\pi$ around an axis $\mathcal{P}$ in the middle of the layers (Fig.3-d): this exchanges the spin batteries and the ferromagnets and the voltage drop gets reversed $V_{c}\longrightarrow-V_{c}$. But one could also rotate the setup by an angle $\pi$ around an axis $\mathcal{F}$ splitting the layers symmetrically (Fig.3-b): all the magnetizations get reversed; the spin current reference axis is also reversed but the charge voltage is obviously unchanged. But a spin current $+I_{s}$ against a given magnetization reference axis $\mathbf{M}$ is equivalent to a spin current $-I_{s}$ with a reversed reference axis $-\mathbf{M}$. One therefore ends up with Fig.3-c which is equivalent to Fig.3-d apart from an opposite voltage. This implies that $V_{c}=0$. The obvious issue with such a setup is that both the spin batteries and the voltage probes must be located at the same positions. This is however easy to relax by doubling the voltage probes and positioning them in a symmetric manner with respect to symmetry plane $\mathcal{F}$ (see Fig.3-e); one would then measure two voltages drops $V_{c1}$ and $V_{c2}$ which might not be offset-free but their average will be: $V_{c}=\left(V_{c1}+V_{c2}\right)/2$.
Figure 3: (a) CPP symmetric spin valve connected to two spin batteries injecting opposite spin currents $I_{s}$ into the ferromagnets. The black arrows show the magnetization direction in each layer ($F1$or $F2$) or the reference direction used to measure the spin current (in the spin batteries). (b) Setup rotated around a symmetry axis $\mathcal{D}$ assumed to exist. (b) Previous setup where the spin currents take opposite values if the reference axis is flipped in the spin batteries. (d) Setup rotated by an angle $\pi$ around a symmetry axis $\mathcal{P}$ assumed to exist in the middle. Comparing (c) and (d) shows they’re identical setups provided the ferromagnets $F1$ $F2$ are identical (for symmetrical voltage probes), so that for anti-parallel magnetization, there is no voltage drop across terminals $T_{1}$ and $T_{2}$. (e) Instead of two voltage probes in the middle of the layers, it might be easier to use 2 sets of voltage probes positioned symmetrically with respect to axis $\mathcal{F}$.
We illustrate such an AP-IMR with an MTJ (magnetic tunnel junction). The junction voltage can be shown using Julliere model [17] and Slonczewski circuit theory [29] to be:
$V_{J}=-\frac{1}{g_{0}}\>\frac{P_{1}+P_{2}}{\left(1-P^{2}\right)^{2}}\>I_{s,J}$ (26)
where $P_{1/2}$ are the usual DOS polarizations, $P=\left|P_{1/2}\right|$, $g_{0}=\Gamma\>n_{1}\>n_{2}$ (with $\Gamma$ the tunneling probability, $n_{1/2}$ the total density of states) and $I_{s,J}$ the spin current across the junction. This obviously vanishes for antiparallel magnetizations which implies an anti-parallel IMR. This calculation neglects however the spin accumulation effects in the ferromagnets. For very thin ferromagnets, a more complete calculation incorporating the spin accumulation in the ferromagnets, leads to
$TMR=\frac{V_{c}^{P}-V_{c}^{AP}}{V_{c}^{AP}}\sim\frac{P_{c}\,r_{c}}{P_{F}\,r_{F}\>\delta l}$ (27)
[…]
Let us consider layers which are not infinitely thin; for instance, we assume equal widths $d_{1}=d_{2}=5\>nm$ up to a monolayer; this means $l_{1/2}=d_{1/2}/l_{F}=0.125$. Using the same figures as above numbers, a more realistic calculation [27] shows that the spin current asymmetry can usually be neglected and leads to $TMR\sim 10^{5}\,\%$, which is still extremely large.
For an FNF trilayer in the same geometry, in the limit of thin layers, one gets a similar MR ratio. The ratio is however much higher for an MTJ because $r_{c}$ is considerably larger by a factor $10^{3}-10^{5}$. For all metallic spin valves, a spin battery based on ferromagnetic resonance would be able to achieve HIGH signals in the $1-100\>\mu V$ range. By using an MTJ, since the resistance is multiplied by a factor $10^{3}-10^{5}$, this implies that $mV$ signals are very achievable which would make such a spin battery driven MTJ quite useable.
What would be the impact of MgO barriers ? This would increase the value of $P_{c}$ and the TMR ratio. But it is not as critical as having very small widths for the ferromagnets as well as a very small width asymmetry. For instance, let us consider figures extracted from Yuasa and coll.[33]. $RA\sim 10-10^{7}\Omega\mu m^{2}$ and the TRM ratio reaches $180\>\%$ so $P_{1}=0.69$ or $P_{c}=0.93$. If we plug this in the previous TMR ratio calculation, this would increase it by a factor 1.5 which is nice but not as crucial as having a very small $\delta l$.
4 Conclusion.
We have introduced several spin valve setups which due to symmetry can exhibit both the bipolar effect of Johnson spin transistor and an IMR. An obvious application of IMR and BE would be more sensitive magnetic sensors so an experimental confirmation of these predictions would be very useful. An essential component for the occurence of both BE and IMR is to replace charge batteries by spin batteries since this allows the occurence in response functions of odd terms $\propto sgn\left(M_{z}\right)$. Since real systems are never perfectly symmetric, we have also quantified the impact of length & spin current asymmetries. The MR ratio can still reach high ranges $10-10^{3}$. The simplest system to observe the BE is the twin system with shared spin battery; for observation of IMR, it is better to test on a twin system with independent spin batteries. The CPP MTJ with two spin batteries is quite promising in comparison with the CPP FNF trilayer due to the much larger HIGH signal which should make the $mV$ range quite reachable. Extensions of this work to non-collinear systems with symmetries, to CIP setups, to parallel IMR in CPP systems and inclusion of non-ideal spin batteries are discussed elsewhere [27].

### v3

@@
 It is closely related to the bipolar effect (BE) of Johnson transistor which can be triggered under the same symmetry conditions.
-1 Introduction
+I Introduction
 This paper introduces a new magneto-resistive effect, the infinite magneto-resistance (IMR) related to the giant magneto-resistance spin valve (GMR [1, 2]) and to the bipolar spin transistor (BST) of Johnson [3, 4].
-The IMR does not rely on specific materials such as half-metals to increase GMR (Heusler alloys as in [5]) or coherent tunneling through MgO barriers [6, 7].
+The IMR does not rely on specific materials such as half-metals to increase GMR (Heusler alloys as in Kimura et al[5]) or coherent tunneling through MgO barriers [6, 7].
 It consists in achieving a zero signal for a system with 2 magnetically active parts (like two ferromagnetic layers in a spin valve) so that the optimistic (resp.
@@
 This is an example of what we will christen as ’twin systems’.
+In Fig.
+1, the ferromagnets can be thought of being in a non-local geometry.
+But they could also be thin films as in Fig.
+2 -a
 Figure 1: (a) The voltage measured across two copies of the same system (and batteries) vanishes by symmetry.
+(b) If one ferromagnet is flipped, one gets a non-vanishing voltage.
+(c) If both ferromagnets are flipped from (a), the voltage still vanishes.
+(d) The voltage is now the opposite of (b) since (d) follows from (b) by flipping both magnetizations, which amount to exchanging both systems.
+Another example of a ’twin system’ is given in Fig.
+2-b.
+It is an $FNF$ valve in a non-local geometry with identical ferromagnets and layers parallel to the figure plane; it could also be considered to be a trilayer with layers perpendicular to the figure plane.
+A spin battery injects a spin current through the intermediate $N$ layer.
+The charge voltage across the trilayer obviously vanishes when the ferromagnets have parallel magnetizations but is non-zero for anti-parallel ones.
+Figure 2: (Color online) (a) The voltage measured across two copies of the same system (and batteries) vanishes by symmetry.
 (b) FNF valve connected to a spin battery and a voltmeter.
@@
 For parallel magnetizations, the voltage again vanishes (for symmetric probes).
+(e) Twin system with an angular offset.
 The black arrows depict the magnetization directions.
-Another example of a ’twin system’ is given in Fig.
-1-b.
-It is an $FNF$ valve in a non-local geometry with identical ferromagnets and layers parallel to the figure plane; it could also be considered to be a trilayer with layers perpendicular to the figure plane.
-A spin battery injects a spin current through the intermediate $N$ layer.
-The charge voltage across the trilayer obviously vanishes when the ferromagnets have parallel magnetizations but is non-zero for anti-parallel ones.
-In practice the vanishing of the signal is constrained by the degree of symmetry.
-Suppose for instance that the widths of the ferromagnets are equal to $l_{m}\pm\delta l$.
-Then we expect a scaling:
+In practice the vanishing of the signal is constrained by the degree of symmetry. Suppose for instance that the widths of the ferromagnets are equal to $l_{m}\pm\delta l$. Then we expect a scaling:
 $\frac{V_{c}^{AP}-V_{c}^{P}}{V_{c}^{P}}\sim\frac{l_{m}}{\delta l}$ (4)
@@
 Fig.
-1-c shows the twin system from Fig.
-1-b with an FNF nanopillar acting as a spin battery.
+2-c shows the twin system from Fig.
+2-b with an FNF nanopillar acting as a spin battery.
 There are other possible setups than ’twin setups’ which will exhibit IMR and bipolarity, as a result of symmetry.
 For instance consider a CPP symmetric trilayer where instead of a charge battery, one connects the layers to two spin batteries as in Fig.
-1-d.
-(The CPP qualifier, current perpendicular to plane, comes from the fact that if one replaced the voltage probes by current ones, the charge current would be perpendicular to the layers.) This setup exhibits IMR for identical incoming spin currents and parallel magnetizations.
-We will discuss later a similar setup with opposite spin currents; one then has IMR for antiparallel magnetizations.
-The latter case is actually easy to understand in terms of a resistor analogy.
-We consider a spin accumulation source for the spin battery (a spin current source will yield similar results); in this example we will measure a charge current.
-In the resistor analogy for a CPP symmetric spin valve with anti-parallel magnetizations driven by a charge battery, currents flowing in each spin channel are equal $I_{\uparrow}=I_{\downarrow}$.
-The key observation is to ask what would happen if the spin channels had a reverse bias (opposite voltages) as in Fig.
-2-(b) : the currents in each spin channel are now opposite $I_{\uparrow}=-I_{\downarrow}$; obviously no charge current can thus flow out of the device (since $I_{c}=I_{\uparrow}+I_{\downarrow}$) in contrast to Fig.2-(a) (parallel case).
-But a reverse bias is precisely what a spin accumulation source can achieve: indeed a spin battery splits the chemical potentials for each spin channel $V_{\uparrow}\neq V_{\downarrow}$ and additionally $V_{\uparrow}=-V_{\downarrow}$ since $V_{c}=0$ in the absence of a charge voltage applied to the system.
-Replacing a charge battery with a spin battery therefore generates reversed biases and opposite currents in each spin channel (for a symmetric spin valve), so that in the anti-parallel state, there is no outgoing charge current: the corresponding ratio $\left(I_{c}^{AP}-I_{c}^{P}\right)/I_{c}^{AP}$ goes to infinity (IMR effect).
-Figure 2: In the resistor network model, symmetric CPP spin valve driven by a spin battery which injects a spin current and induces spin accumulation.
-The latter is equivalent to reverse biasing the two spin channels.
-The currents (in electrical units) in each spin channel are $I_{\uparrow}$ and $I_{\downarrow}$.
-(a) Parallel magnetizations for ferromagnets $F1$ and $F2$ : non-vanishing charge current $I_{c}^{P}\neq 0$.
-(b) Anti-parallel magnetizations: current $I_{c}$ at detector is vanishing since the currents in each channel are opposite.
-Therefore one has an infinite resistance state.
-2 Response function, BE & IMR.
-2.1 Trans-resistance.
+2-d.
+(By CPP, current perpendicular to plane, we mean that if one replaced the voltage probes by current probes, the charge current would flow perpendicular to the layers.) This setup exhibits IMR for identical incoming spin currents and parallel magnetizations (parallel IMR).
+We will discuss later a similar setup with opposite spin currents; one then has IMR for antiparallel magnetizations (AP IMR).
+Note also that, as shown on 2-e, the two copies of the same system can have arbitrary relative orientations provided the voltage measurement probes remain symmetric.
+Figure 3: (Color online) Resistor model of a spin valve driven by a spin current source.
+The charge voltage is $V_{c}=\left(V_{\uparrow}+V_{\downarrow}\right)/2$.
+(a) $V_{c}\neq 0$ when ferromagnets $F1$ and $F2$ have parallel magnetizations.
+(b) But for anti-parallel magnetizations, the voltage $V_{c}$ obviously vanishes since $V_{\uparrow}=\left(r+R\right)\>I_{s}/2=-V_{\downarrow}$.
+The AP IMR for CPP spin valve is actually easy to understand in terms of a resistor analogy with a single spin battery (due to current conservation on each spin channel, a second spin battery is irrelevant, see Fig.4-a).
+We consider a spin current source in a two-channel model ( Fig.
+3) : this is tantamount to having opposite charge current sources for each spin channel.
+In the resistor analogy for a CPP symmetric spin valve with anti-parallel magnetizations driven by a charge battery, the currents flowing in each spin channel would be equal $I_{\uparrow}=I_{\downarrow}$ and the voltages for each spin channel would be identical in the anti-parallel case $V_{\uparrow}=V_{\downarrow}$ In the antiparallel case for a spin current source driving the resistors ( Fig.3-(b) ), since the currents are opposite $I_{\uparrow}=-I_{\downarrow}=I_{s}/2$, the channel voltages are also opposite $V_{\uparrow}=-V_{\downarrow}$ so that $V_{c}=0$.
+The charge voltage divided by the spin current has the dimension of a resistance $R=V_{c}/I_{s}$.
+The corresponding MR ratio $\left(R^{AP}-R^{P}\right)/R^{AP}$ goes to infinity (IMR effect).
+A similar resistor analogy can be drawn for parallel IMR for CPP spin valves but requires to take into account spin flips and transfer from one spin channel to the other.
+See Fig.4-b.
+Figure 4: (Color online) (a) Resistor analogy for AP IMR of CPP spin valve connected to two spin batteries.
+Since spin flip is neglected, the spin current source on the right can be removed because it is redundant.
+(b) The resistor analogy for parallel IMR of CPP spin valve connected to 2 spin current sources requires addition of spin mixing either in bulk and/or at the interface between the ferromagnets.
+In the resistor circuit, the resistance $r^{\prime}$ ensures the current to be in a closed loop while no current flows through resistances $r$ and $R$.
+Therefore the charge voltage $V_{c}=\left(V_{\uparrow}+V_{\downarrow}\right)/2$ vanishes.
+The same conclusion is reached by adding a mixing resistance at the interface between the ferromagnets.
+II Response function, BE & IMR.
+II.1 Trans-resistance.
 Let us define a charge-spin trans-resistance encoding the response to a spin battery:
-$r_{cs}=\left(\frac{\partial V_{c}}{\partial I_{s}}\right)_{I_{c}=0}.$ (5)
+$r_{cs}=\left(\frac{\partial V_{c}}{\partial I_{s}}\right)_{I_{c}=0},$ (5)
+where the spin current is defined in electrical units as $I_{s}=I_{\uparrow}-I_{\downarrow}$ for collinear setups.
 It has the dimension of a resistance.
@@
 The resistance $r_{cs}$ can take four independant values.
-2.2 Bipolar effect (BE).
+II.2 Bipolar effect (BE).
 This requires the concurrent switching of both magnetizations. One has a bipolar effect if for some values of $\sigma_{1}$ and $\sigma_{2}$
@@
 For Fig.
-1-b, eq.(8) is also true since a symmetry through plane $(\mathcal{P})$ parallel to the layers and perpendicular to the Figure maps each ferromagnetic layer unto the other but leaves the spin battery invariant.
+2-b, eq.(8) is also true since a symmetry through plane $(\mathcal{P})$ parallel to the layers and perpendicular to the Figure maps each ferromagnetic layer unto the other but leaves the spin battery invariant.
 But the voltage probes are exchanged across that plane so that $V_{c}\longrightarrow-V_{c}$.
@@
 $r_{cs}\left(\sigma_{1},\,-\sigma_{1}\right)=2r_{1}\sigma_{1}=-r_{cs}\left(-\sigma_{1},\,\sigma_{1}\right).$ (9)
-2.3 IMR.
+II.3 IMR.
 For the twin system of Fig.2-a & b, it also follows from eq.(8) that $r_{cs}^{P}=r_{cs}\left(\sigma_{1},\,\sigma_{1}\right)=0$ while $r_{cs}^{AP}\neq 0$.
@@
 Since the condition for IMR also allows the BE, let us compare the two effects for the setups of Fig.
-1-a, 1-b.
+2-a, 2-b.
 IMR is observed by switching a single magnetization while bipolarity requires switching two magnetizations, which may or may not be possible or desireable in a given setup.
@@
 The switching field needs to be tailored according to the effect one wishes to trigger, whether it be IMR or the bipolar effect.
-3 Example setups.
-3.1 Twin setup with shared spin battery.
+III Example setups.
+III.1 Twin setup with shared spin battery.
 To quantify the impact of asymmetries, one can make explicit calculations for the geometry of Fig.(2-b) in the one dimensional limit, using Valet-Fert drift diffusion model.
@@
 one finds after tedious calculations that $r_{cs}$ splits into 3 parts, a bulk contribution (I) and two interfacial ones (II & III):
-$r_{cs}^{I}$ $=\left(P_{F1}-P_{F2}\right)\>\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right]\>\frac{r_{F}}{2}+\delta l\>\left(P_{F1}+P_{F2}\right)\>\frac{r_{F}}{2}\>\alpha,$ (11)
+$r_{cs}^{I}$ $=\left(P_{F1}-P_{F2}\right)\>\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right]\>\frac{r_{F}}{2}$
+$+\delta l\>\left(P_{F1}+P_{F2}\right)\>\frac{r_{F}}{2}\>\alpha,$ (11)
 $r_{cs}^{II}$ $=\left(P_{c1}-P_{c2}\right)\>\frac{r_{c}}{2}+\delta l\>\left(P_{c1}+P_{c2}\right)\>\frac{r_{c}}{2}\>\beta,$ (12)
-where $\alpha$ and $\beta$ are factors which depend on the geometry and the various spin resistances (for their exact expressions, we refer to [27]).
-$r_{cs}^{III}$ $=\left(P_{c1}^{\prime}-P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)+\delta l\>\left(P_{c1}^{\prime}+P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$ (13)
+where $\alpha$ and $\beta$ are factors which depend on the geometry and the various spin resistances, namely
+$\alpha$ $=\left[\cosh_{m}-\left(f_{m}^{2}-1\right)\left(\cosh_{m}-1\right)-f_{m}\sinh_{m}\right]$
+$+\frac{f_{m}^{2}-1}{\frac{r_{c}}{r_{F}}+f_{m}}\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right],$ (13)
+$\beta$ $=\frac{f_{m}^{2}-1}{\frac{r_{c}}{r_{F}}+f_{m}}.$ (14)
+$r_{cs}^{III}$ $=\left(P_{c1}^{\prime}-P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$ (15)
+$+\delta l\>\left(P_{c1}^{\prime}+P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$
 One can compute a resistance ratio as $MR=\frac{R_{AP}-R_{P}}{R_{P}}$ where the subscripts $P/AP$ refer to parallel or anti-parallel magnetizations; it follows that:
@@
 If the widths are about $10\pm 0.5\,nm$, $\delta l\sim\frac{1}{60}$ so up to numerical factors $MR\sim 6000\%.$
-It is also important to have an idea of the range of the HIGH signal.
+It is also important to have an idea of the range of the HIGH signal (the LOW signal being close to 0).
 Let’s use MSU results for CoCu at $4.2\>K$ and for simplicity assume that $P_{c}\sim P_{c}^{\prime}=0.77\pm 0.05$, $P_{F}=0.46\pm 0.05,$ $2r_{c}A\sim 2r_{c}^{\prime}A=1.02\pm 0.04\>f\Omega m^{2}$, $t_{Co}=2-10\,nm,\>l_{Co}>60\>nm$ and $\rho_{F}(Co)=75\pm 5\>n\Omega m,$ $\rho_{N}(Cu)=6\pm 1\>n\Omega m.$ Then $X\sim 1$, $f_{m}\sim 1$.
@@
 This will be in the range of $R=1\>\Omega$ for $A=10^{-3}\mu m^{2}$ which corresponds to a nanopillar.
-3.2 Twins with individual spin batteries.
+III.2 Twins with individual spin batteries.
 We consider two identical systems connected to a spin battery and interconnected by voltage probes (Fig.(2-a)).
@@
 The HIGH signal is dominated by interfaces in the limit of very thin ferromagnets: $r_{cs}^{high}\longrightarrow 2\>\left\{P_{c}\,r_{c}+P_{c}^{\prime}\,r_{c}^{\prime}\right\}$ while for very thick layers: $r_{cs}^{high}\longrightarrow 2\>\left\{P_{c}\,r_{c}+P_{F}\,r_{F}\right\}.$ Using the same parameters as above, one finds again a range $r_{cs}^{high}\,A\sim 1\>f\Omega m^{2}.$
-3.3 Symmetric CPP spin valves operated by two spin batteries. Parallel IMR.
+III.3 Symmetric CPP spin valves operated by two spin batteries. Parallel IMR.
 One can define new transport coefficients per:
@@
 Note that the setup in Fig.
-1-d could represent also a nano-pillar or a non-local geometry in addition to the usual multi-layer thin films.
+2-d could represent also a nano-pillar or a non-local geometry in addition to the usual multi-layer thin films.
 The spin current from the batteries are oriented to be incoming in the load (outgoing from the batteries).
@@
 If one starts with antiparallel magnetizations and switches both of them, one gets a bipolar effect.
+See Fig.4 for the resistor analogy.
 After rotating the setup around an axis in the middle of the paramagnetic layer:
@@
 $V_{c}\left(\sigma_{1},\>\sigma_{1}\right)=-V_{c}\left(\sigma_{1},\>\sigma_{1}\right)=0$ (26)
-and a bipolar effect for antiparallel magnetizations when flipping both of them for identical currents: $V_{c}\left(\sigma_{1},\>-\sigma_{1}\right)=-V_{c}\left(-\sigma_{1},\>\sigma_{1}\right).$ Detailed calculations on an FNF trilayer exhibiting such a parallel IMR are given elsewhere [27].
-In the limit of infinite ferromagnets, one finds
+and a bipolar effect for antiparallel magnetizations when flipping both of them for identical currents: $V_{c}\left(\sigma_{1},\>-\sigma_{1}\right)=-V_{c}\left(-\sigma_{1},\>\sigma_{1}\right).$ Detailed calculations on an FNF trilayer exhibiting such a parallel IMR will be given elsewhere.
+In the limit of infinite ferromagnets, one finds in terms of a mean spin current $I_{s,0}^{m}=\left(I_{s,L}+I_{s,R}\right)/2$ and a spin current asymmetry $\delta I_{s}=\left(I_{s,L}-I_{s,R}\right)/2$:
 $MR=\frac{V_{c}^{AP}-V_{c}^{P}}{V_{c}^{P}}\sim\frac{I_{s,0}^{m}}{\delta I_{s}}$ (27)
-3.4 Anti-parallel IMR in CPP MTJ & FNF trilayer.
-It is also possible to engineer an anti-parallel IMR.
+(Length asymmetry drops out for infinite systems.)
+III.4 Anti-parallel IMR in CPP MTJ & FNF trilayer.
+It is also possible to engineer an anti-parallel IMR (see Fig.3 for the resistor analogy explaining it).
 Fig.5-a shows such a setup.
@@
 This is however easy to relax by doubling the voltage probes and positioning them in a symmetric manner with respect to symmetry plane $\mathcal{F}$ (see Fig.5-e); one would then measure two voltages drops $V_{c1}$ and $V_{c2}$ which might not be offset-free but their average will be: $V_{c}=\left(V_{c1}+V_{c2}\right)/2$.
-Figure 3: (a) CPP symmetric spin valve connected to two spin batteries injecting opposite spin currents $I_{s}$ into the ferromagnets.
+Figure 5: (Color online) (a) CPP symmetric spin valve connected to two spin batteries injecting opposite spin currents $I_{s}$ into the ferromagnets.
 The black arrows show the magnetization direction in each layer ($F1$or $F2$) or the reference direction used to measure the spin current (in the spin batteries).
@@
 (b) Previous setup where the spin currents take opposite values if the reference axis is flipped in the spin batteries.
-(d) Setup rotated by an angle $\pi$ around a symmetry axis $\mathcal{P}$ assumed to exist in the middle.
+(c) Setup rotated by an angle $\pi$ around a symmetry axis $\mathcal{P}$ assumed to exist in the middle.
+(d) is (b) with reference directions flipped in the spin batteries (which changes the sign of spin currents).
 Comparing (c) and (d) shows they’re identical setups provided the ferromagnets $F1$ $F2$ are identical (for symmetrical voltage probes), so that for anti-parallel magnetization, there is no voltage drop across terminals $T_{1}$ and $T_{2}$.
 (e) Instead of two voltage probes in the middle of the layers, it might be easier to use 2 sets of voltage probes positioned symmetrically with respect to axis $\mathcal{F}$.
+The previous discussion then applies to $V_{c}=\left(V_{c1}+V_{c2}\right)/2$.
 We illustrate such an AP-IMR with an MTJ (magnetic tunnel junction). The junction voltage can be shown using Julliere model [17] and Slonczewski circuit theory [28] to be:
@@
 Let us consider layers which are not infinitely thin; for instance, we assume equal widths $d_{1}=d_{2}=5\>nm$ up to a monolayer; this means $l_{1/2}=d_{1/2}/l_{F}=0.125$.
-Using the same figures as above numbers, a more realistic calculation [27] shows that the spin current asymmetry can usually be neglected and leads to $TMR\sim 10^{5}\,\%$, which is still extremely large.
+Using the same figures as above numbers, a more realistic calculation (which will be detailed elsewhere) shows that the spin current asymmetry can usually be neglected and leads to $TMR\sim 10^{5}\,\%$, which is still extremely large.
 For an FNF trilayer in the same geometry, in the limit of thin layers, one gets a similar MR ratio.
-The ratio is however much higher for an MTJ because $r_{c}$ is considerably larger by a factor $10^{3}-10^{5}$.
-For all metallic spin valves, a spin battery based on ferromagnetic resonance would be able to achieve HIGH signals in the $1-100\>\mu V$ range.
+The ratio is however much higher for an MTJ because $r_{c}$ is considerably larger, by a factor $10^{3}-10^{5}$.
+For all metallic spin valves, a spin battery based on ferromagnetic resonance would be able to achieve signals in the $1-100\>\mu V$ range.
 By using an MTJ, since the resistance is multiplied by a factor $10^{3}-10^{5}$, this implies that $mV$ signals are very achievable which would make such a spin battery driven MTJ quite useable.
@@
 If we plug this in the previous TMR ratio calculation, this would increase it by a factor 1.5 which is nice but not as crucial as having a very small $\delta l$.
-4 Conclusion.
+IV Conclusion.
 We have introduced several spin valve setups which due to symmetry can exhibit both the bipolar effect of Johnson spin transistor and an IMR.
@@
 The CPP MTJ with two spin batteries is quite promising in comparison with the CPP FNF trilayer due to the much larger HIGH signal which should make the $mV$ range quite reachable.
-Extensions of this work to non-collinear systems with symmetries, to CIP setups, to parallel IMR in CPP systems and inclusion of non-ideal spin batteries are discussed elsewhere [27].
+Extensions of this work to non-collinear systems with symmetries, to CIP setups, to parallel IMR in CPP systems and inclusion of non-ideal spin batteries are discussed elsewhere.